OpenStudy (anonymous):

If sin 2 theta = cos 3, then theta is equal to ( theta<= 90°)

OpenStudy (anonymous):

\[\sin 2 \theta=2 \sin \theta \cos \theta\]

OpenStudy (anonymous):

2 sin theta cos theta=cos 3

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

do u have options ?

ganeshie8 (ganeshie8):

from the graph, there are two solutions. not sure how to find them analytically

OpenStudy (anonymous):

18,30 i think

OpenStudy (anonymous):

18,36,18 or 90

OpenStudy (anonymous):

wait cos 3 ? or cos 3 theta ?

OpenStudy (anonymous):

3 degree or rad ?

OpenStudy (anonymous):

anyway , theta = sin^-1 (cos 3 )/2

OpenStudy (anonymous):

your not allowed to use calculator ?

OpenStudy (anonymous):

cos 3

OpenStudy (anonymous):

:O and no calculator ?

OpenStudy (anonymous):

just check options if u have ;)

OpenStudy (anonymous):

theta = sin^-1 (cos 3 )/2 how to get the answer for sin^-1